The magnitude and directions of u and v are given. Find vector w's polar coordinates.
u: magnitde 23, bearing 215°
v: magnitude 14.5, bearing 105°
w: u + v
A:(22.7, 77°)
B:(77, 22.7°)
C:(22.6, 177.9°)
D:(177.9, 22.6°)

Answer by stanbon(75887)

The magnitude and directions of u and v are given. Find vector w's polar coordinates.
u: magnitude 23, bearing 215°
v: magnitude 14.5, bearing 105°
w: u + v
---
x ccordinate of u+v: 23cos(215)+14.5cos(105) = -22.59
y coordinate of u+v: 23sin(s15)+14.5sin(105) = 0.8136
---
magnitude of u+v: sqrt[(-22.59)^2+(0.8136)^2] = 22.60
direction of u+v: tan^-1(0.8136/(-22.59)) = -2.06 = 180-2.06 = 177.9 degrees
----
Ans: C
==================
Cheers,
Stan H.
============================
A:(22.7, 77°)
B:(77, 22.7°)
C:(22.6, 177.9°)
D:(177.9, 22.6°)